Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 10 January
to 6 pm 11 January 1998

Activity at the volcano today was similar to that experienced yesterday although there was a slight increase in the number of rockfall events. Between 16:00 yesterday and 16:00 today there were 18 hybrid earthquakes, 29 rockfall signals, 8 long-period earthquake and 2 volcano-tectonic earthquakes. One of the hybrid earthquakes and 5 of the long period earthquakes preceded rockfall signals. The 6 to 8 hour cyclicity of recent days has continued today with a further reduction in the amplitude of the peaks.

Cloud continues to mask the upper parts of the dome and did not permit any views of the summit. However, seismic rockfall activity suggests continued dome growth in the Galway's sector.

A GPS survey of the network on the north-western flanks of the volcano including Old Towne, Waterworks, Garibaldi, Lees and St. George's Hill was completed today. The results will be processed and reported later this week.

Heavy rain this afternoon resulted in hot mudflows being generated in the Belham River valley. Large logs up to 6 m long and rocks up to 60 cm diameter were transported with the muddy water which was about 10 m wide at Belham bridge. Temperatures of the water about one hour after the peak in activity gave temperatures of only 27.4 degrees centigrade (approximately 9 degrees above air temperature), but rocks out of the main flow on the golf course were still steaming and were hot to touch. Erosion of the recent deposits by the torrents near Molyneaux produced small ash clouds as the sides of the river tumbled into the water.

Dust levels at all monitoring stations across the island were low today.

People should remain alert and listen to Radio Montserrat for further information.

Montserrat Volcano Observatory